OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (2, 3). x = 2 x = 3 y = 2 y = 3

OpenStudy (ledah):

have you tried drawing the line out?

OpenStudy (anonymous):

no can u teach me how to do tat

OpenStudy (ledah):

Absolutely

OpenStudy (anonymous):

thankz

OpenStudy (anonymous):

it's x = 2

Parth (parthkohli):

You got a better way than drawing it out :)

OpenStudy (anonymous):

how

OpenStudy (ledah):

how do answers help him moha? :l

Parth (parthkohli):

\( \color{Black}{\Rightarrow y = 0x + b}\) See how it is a 0 slope?

Parth (parthkohli):

\( \color{Black}{\Rightarrow y = b}\) Now just find the y-intercept.\(\)

OpenStudy (ledah):

Well that way works too. but I learned by drawing it out.

OpenStudy (anonymous):

]Identify the correct x intercept and y intercept of the equation 2x - 3y = -18. x intercept = (9, 0); y intercept = (0, -6) x intercept = (6, 0); y intercept = (0, -9) x intercept = (-9, 0); y intercept = (0, 6) x intercept = (-6, 0); y intercept = (0, 9)

OpenStudy (anonymous):

im so bad at this subject

OpenStudy (anonymous):

sorry i was wrong

OpenStudy (anonymous):

theres so much to no

Parth (parthkohli):

Well, as y doesn't change, then we have y intercept as \((0,2)\)

OpenStudy (ledah):

Math is not my best subject either

Parth (parthkohli):

So, \(y = b = 2\) \(y = 2\)

OpenStudy (ledah):

but I'm great at some parts in it

OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation 2x - 3y = -18. x intercept = (9, 0); y intercept = (0, -6) x intercept = (6, 0); y intercept = (0, -9) x intercept = (-9, 0); y intercept = (0, 6) x intercept = (-6, 0); y intercept = (0, 9) can u help on this

OpenStudy (anonymous):

parth u r greate

Parth (parthkohli):

@zahabaha Can you post it one at a time? Thanks!

Parth (parthkohli):

Thank you @moha_10 :)

OpenStudy (anonymous):

math is my best subject but its just that i havent learned thiis stuff yet

Parth (parthkohli):

Same here lol

OpenStudy (anonymous):

:0